OpenStudy (anonymous):

compute the critical value z a/2 that corresponds to a 93% level of confidence.

OpenStudy (anonymous):

You want to find \(Z_{\alpha/2}\) such that \[P(Z>Z_{\alpha/2})=\frac{1-0.93}{2}=0.035\] assuming \(Z_{\alpha/2}>0\). Graphically, you want to find the dotted lines that separate the population into chunks of \(0.93\) and \(\dfrac{1-0.93}{2}\):  This is a matter of referring to a table of \(z\) values. The probability above is a right-tail probability, so you have to make sure you use the correctly-tailed table of values. Here I'll link you to a left-tail table: http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf In this case, we have the equivalence \[P(Z>Z_{\alpha/2})=P(Z<-Z_{\alpha/2})=0.035\] We find that \(-Z_{\alpha/2}\) must be about \(-2.70\), so \(Z_{\alpha/2}=2.70\).